175x + 200y = 7800
solve for x

Answer by mananth(16946)

175x + 200y = 7800
divide by 25
7x+8y=312
add -8y
7x+8y-8y=312-8y
7x=312-8y
/7
x= 312/7 -8y/7
...
m.ananth@hotmail.ca